OpenStudy (anonymous):

Solve 3x - 9 + x > 27

HanAkoSolo (jamierox4ev3r):

to solve problems like these, generally you have to add the like terms together first. in this case, add 3x and x together. Then, you get 4x. So your problem looks like this thus far: 4x-9>27 then, add 9 to both sides 4x-9>27 +9 +9 -------- 4x>36 divide both sides by 4 4x>36 /4 /4 x=? So what's the answer? @IhateMathDotCom

HanAkoSolo (jamierox4ev3r):

u there? O_O

OpenStudy (anonymous):

Yeah e.e But I honestly just stopped reading at ''add 9 to both sides'' because like where do I add it? Under which 2 numbers?

OpenStudy (anonymous):

Oh wait, nevermind lol thanks

HanAkoSolo (jamierox4ev3r):

got it? :P lol so erm do you get it?

OpenStudy (anonymous):

4/4 = 1 36/4 = 8

OpenStudy (anonymous):

So it's one? Or 9.

HanAkoSolo (jamierox4ev3r):

36/4 actually equals 9, so that means x=9. That's the answer to the problem and yes, lol looked like you said 8

OpenStudy (anonymous):

Okay thanks :D Because my teacher never teaches us this in class. And when she does, she stops to insult someone and create a reason to leave class, leaving us confused.

HanAkoSolo (jamierox4ev3r):

really? :o aww that's terrible! Glad I could help :)

OpenStudy (anonymous):

x < 9, correct? Because 9 is greater than 1?

HanAkoSolo (jamierox4ev3r):

yeah!! lol oops >.<

OpenStudy (anonymous):

Thanks :)